OpenStudy (anonymous):

Find the missing value to the nearest hundredth. sin ___ = 7/12 (1 point) 35.69° 30.26° 80.16° 54.31°

OpenStudy (hoblos):

sin x = 7/12 x = arc sin (7/12) = 35.69°

OpenStudy (anonymous):

35.69°